The question is 2 times the square root of 32 PLUS 3 times the square root of 8 MINUS 4 times the square root of 18.
Please help.

Answer by nyc_function(2741)

Let sqrt = square root for short


sqrt[32] + 3(sqrt[8]) - 4(sqrt[18]) =


Break down every square root. Your goal is to get the same radicand for every term.

sqrt[32] = 8(sqrt[2])


3(sqrt[8]) = 6(sqrt[2])


4(sqrt[18]) = 12(sqrt[2])


We now have 8(sqrt[2]) + 6(sqrt[2]) - 12(sqrt[2])


Notice how every radical has the same number 2. This was the goal.


We now simplify 8(sqrt[2]) + 6(sqrt[2]) - 12(sqrt[2]).


8(sqrt[2]) + 6(sqrt[2]) - 12(sqrt[2]) = 14(sqrt[2]) - 12(sqrt[2]).



14(sqrt[2]) - 12(sqrt[2]) = 2(sqrt[2]) as the final answer.


Did you follow?